Suppose the house costs $200,000. You need to borrow $180,000 from the bank that you will pay over 30 years. Assume the bank charges 3% APR on your mortgage and payments are made at the end of each month. Calculate your monthly mortgage payment. Briefly show your calculations.
Answer #1
Monthly mortgage payment = Loan amount / Present value of annuity of 1
= $       1,80,000 / 237.1894
= $           758.89
Working:
Present value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.0025)^-360)/0.0025 i = 3%/12 = 0.0025
= 237.1893815 n = 30*12 = 360